SOLVE BY COMPETING THE SQUARE:
x^2-6x+12=0
PLEASE HELP... I KNOW HOW TO USE THE QUADRATIC FORMULA... BUT I DO NOT KNOW HOW TO DO THE COMPLETING THE SQUARE...

Answer by checkley77(12569)

x^2-6x+12=0
FIRST REMIOVE THE 12 OVER TO THE RIGHT SIDE OF THE = SIGN.
X^2-6X=12
NOW CALCULATE THE NUMBER TO COMPLETE THE SQUARE.
(6/2)^2=3^2=9 WHICH IS THE MISSING TERM THAT MUST BE ADDED TO EACH SIDE OF THE EQUATION
X^2-6X+9=12+9
(X-3)^2=21
TAKE THE SQUARE ROOT OF BOTH SIDES
X-3=SQRT21
ADD 3 TO BOTH SIDES
X=SQRT21+3
X=4.58+3
X=7.58 ANSWER.